# Whom We Trust, What We Fear: COVID-19 Fear and the Politics of Information

### v1

@@ Abstract.
The COVID-19 pandemic triggered not only a global health crisis but also an infodemic, an overload of information from diverse sources influencing public perception and emotional responses. In this context, fear emerged as a central emotional reaction, shaped by both media exposure and demographic factors. In this study, we analyzed the relationship between individuals’ self-reported levels of fear about COVID-19 and the information sources they rely on, across nine source categories, including medical experts, government institutions, media, and personal networks. In particular, we defined a score that ranks fear levels based on self-reported concerns about the pandemic, collected through the Delphi CTIS survey in the United States between May 2021 and June 2022. We found that both fear levels and information source usage closely follow COVID-19 infection trends, exhibit strong correlations within each group (fear levels across sources are strongly correlated, as are patterns of source usage), and vary significantly across demographic groups, particularly by age and education. Applying causal inference methods, we showed that the type of information source significantly affects individuals’ fear levels. Furthermore, we demonstrated that information source preferences can reliably match the political orientation of U.S. states. These findings highlight the importance of information ecosystem dynamics in shaping emotional and behavioral responses during large-scale crises.
Fear, Source of information, COVID-19, Causal inference, Online survey, United States
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: 32st SIGKDD Conference on Knowledge Discovery and Data Mining; August 09–13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2018/06ccs: Mathematics of computing Causal networksccs: Applied computing Health informatics
1. Introduction
@@
The COVID-19 pandemic represented a global health crisis of unprecedented scale and complexity. At the same time, it generated an infodemic—a rapid and overwhelming surge of information from a wide range of sources (Solomon et al., 2020). As individuals sought clarity and guidance, they turned to news media, social platforms, government announcements, and public health organizations. The credibility, consistency, and framing of these sources played a critical role in shaping public understanding of the threat. Additionally, they influenced emotional responses across populations (Jones et al., 2021; Lu et al., 2021).
@@
In this context, fear emerged as one of the most prominent psychological responses to the pandemic (Coelho et al., 2020). This emotion was not driven solely by the objective health threat, but also by the volume, tone, and perceived trustworthiness of the information consumed. The impact of these information sources on fear varied across different demographics. Differences in age, education, gender, and socioeconomic status significantly shaped how individuals selected, interpreted, and trusted various sources (Nino et al., 2021). These disparities highlight the necessity of accounting for both the nature of the information disseminated and the characteristics of the recipients when evaluating the psychological impact of information exposure.
@@
Figure 1. Schematic representation of our study.
1.1. Contributions
@@
Our main contributions are as follows—summarized in Figure 1. First, using statistical tests and correlation analyses on U.S. data, we quantify how different sources of information influence individuals’ levels of fear related to COVID-19. Second, we investigate how demographic factors—specifically age and education—not only affect emotional responses but also significantly shape information-seeking behavior and trust, revealing notable patterns across population groups. Third, through a causal inference model (Pearl, 2010), we compare the relative impact of age, education, and information source on fear, finding that information source is the most influential driver. Finally, building on this, we apply K-means clustering (Hartigan et al., 1979) to group U.S. states based on their residents’ reliance on different types of information. Remarkably, the resulting clusters exhibit a strong alignment with state-level political orientation—specifically, whether a state tends to lean Democratic or Republican. Collectively, our findings underscore the critical role of information sources in shaping public responses to large-scale crises. They also highlight the dangers of misinformation—particularly when disseminated by institutional or authoritative entities—which can distort public perception, exacerbate fear, and undermine trust in democratic institutions.
@@
The remainder of this article is organized as follows. Section 2 reviews prior research on fear related to COVID-19, identifying key findings and existing gaps. Section 3 describes the data sources and the construction of the indices used in our analysis. In Section 4, we examine how fear levels vary according to the information sources individuals rely on. Section 5 introduces a causal inference model to identify the variables most strongly associated with fear. In Section 6, we analyze patterns of information reliance across the U.S. population and apply clustering techniques to group states accordingly. Finally, Section 7 summarizes our key findings and discusses their broader implications.
2. Related work
A growing body of research—predominantly based on survey data and often employing standardized tools such as the Fear of COVID-19 Scale (FCV-19S)—has explored the psychological impact of the COVID-19 pandemic, particularly the prevalence and determinants of fear. Developed by Ahorsu et al. (Ahorsu et al., 2022) in Iran, the FCV-19S has enabled researchers to quantify and compare fear levels across populations and has since been validated in various countries, including Spain (Mart´ınez-Lorca et al., 2020), Turkey (Gokseven et al., 2022), and India (Doshi et al., 2021), proving to be a reliable and consistent measure.
Several studies highlight how fear was shaped by both media exposure and interpersonal communication. For instance, Ruobing (Li, 2022) found that heightened fear was linked to increased media use and more danger-focused conversations, creating a feedback loop that sustained anxiety. However, individuals also employed strategies such as avoidance, counterarguing, and reappraisal, with counterarguing emerging as an effective method to reduce fear and encourage continued engagement. Complementing this, a study in Lebanon demonstrated that increased media exposure was positively associated with adherence to preventive behaviors, mediated by perceived knowledge and fear, emphasizing the complex role of media in shaping health-related actions during the pandemic (Melki et al., 2022).
Gender, age, health status, and socio-economic factors also played significant roles in shaping fear responses. Studies from diverse regions—including Chile (Cerda and Garc´ıa, 2022), Turkey (Gokseven et al., 2022), and India (Doshi et al., 2021)—consistently found that women, individuals with chronic illnesses, those with limited knowledge of the pandemic, and people facing economic hardship or living alone experienced greater levels of fear. Conversely, older age, higher education, and greater COVID-19 knowledge appeared to mitigate these fears. These findings underscore the importance of context-specific mental health strategies, particularly for vulnerable demographic groups. Notably, average fear scores varied significantly across population subgroups. For example, Doshi et al. (Doshi et al., 2021) found that while more than half of participants in India reported low levels of fear, certain groups—including women, married individuals, and healthcare workers—exhibited significantly higher fear levels.
Broader reviews and national surveys further enrich this picture. Quadros et al. (Quadros et al., 2021), in a scoping review of early pandemic studies, reported the FCV-19S ranging from 18.1% to 45.2%, particularly among women, healthcare workers, and those with pre-existing mental health concerns. In the U.S., Warren et al. (Warren et al., 2021) highlighted the role of anxiety sensitivity—especially regarding physical symptoms—in driving COVID-19-related fear, which in turn was linked to heightened anxiety and depression. Adding a longitudinal perspective, a 14-month study tracked fear trajectories over time, finding that while fear can be protective by promoting preventive behaviors, excessive fear contributes to anxiety and depression, illustrating the delicate balance between adaptive and maladaptive emotional responses during the pandemic (Mertens et al., 2023).
Overall, these findings emphasize that fear of COVID-19 was not uniformly experienced but varied significantly by individual characteristics and broader contextual factors. As Bhattacharjee et al. (Bhattacharjee and Acharya, 2020) noted, the pandemic’s mental health toll was widespread, with many Americans expressing fear of infection, death, and economic loss early on. This collective distress highlights the urgent need for culturally sensitive, targeted mental health interventions, particularly during public health crises.
@@ 2. Related work
(I5) In the past 7 days, from which of the following sources have you received news and information about COVID-19? (1) Doctors, (2) Scientists, (3) CDC, (4) Government, (5) Politicians, (6) Journalists, (7) Friends and family, (8) Religious leaders, (9) None of the above (other sources not listed here).
@@
Table 1. Selected questions and responses from the Delphi US CTIS survey (Delphi Group, CMU, 2020), with labels as in the original survey.
To summarize, numerous studies have utilized the FCV-19S to examine pandemic-related fear, consistently identifying factors such as media exposure, anxiety sensitivity, and demographic variables—including age and gender—as key predictors. These investigations reveal robust associations between fear, anxiety, and depression across diverse populations, including older adults, healthcare workers, and students. Systematic reviews further highlight the high prevalence of fear during the pandemic and underscore the critical need for targeted mental health interventions.
@@
However, most prior research has been cross-sectional and largely focused on isolated individual predictors, with limited attention to the combined effects of demographics and information sources on fear. Additionally, geographic and political influences remain underexplored. Our study addresses these gaps by employing advanced statistical and causal inference methods to evaluate how demographics and sources of information jointly shape fear responses. Moreover, by clustering U.S. states according to patterns of information source usage and linking these clusters to political orientation, we offer a comprehensive, integrated perspective on the social and contextual determinants of COVID-19 fear.
It is important to note that the FCV-19S was not employed in this study, as it is based on a specific survey (Ahorsu et al., 2022). Instead, as outlined in Section 3, our analysis is based on a distinct dataset collected through an alternative survey. Consequently, direct comparisons with results derived from the FCV-19S are not feasible.
3. Data
@@
Throughout this work, we relied on multiple sources of COVID-19 data, including official surveillance data from the COVID-19 Data Hub (Guidotti and Ardia, 2020; Guidotti, 2022) and survey-based data from the Delphi US CTIS (Delphi Group, CMU, 2020), which contains information on demographics, the types of information sources used by individuals, and self-reported levels of fear. All datasets used in this study are open-source and publicly available. Aggregated data from the Delphi US CTIS survey can be accessed as described in the Availability of materials and data section.
@@
We primarily used the COVID-19 Data Hub (Guidotti and Ardia, 2020; Guidotti, 2022) as our source of surveillance data. Specifically, we used daily new fatalities and confirmed cases—computed from cumulative counts—as well as the total U.S. population. These indicators were used to estimate the daily number of official infections via Sybil (Baccega et al., 2024). Specifically, this data allows us to reconstruct the dynamics of a Susceptible-Infected-Recovered-Deceased-Susceptible (SIRDS) compartmental model with reinfections over the study period. In this model, the infected compartment represents the estimate of the official number of infections—we assumed an average recovery time of 14 days to compensate for the lack of recovery data, which is largely unavailable. While the Data Hub provides additional information, such as vaccination rates and hospital statistics, these were not relevant to our study.
@@
Since the spring of 2020, Facebook has collaborated with both Carnegie Mellon University (CMU) and the University of Maryland (UMD) through its Data for Good program (Facebook, 2020) to promote large-scale, daily surveys developed by these institutions (Astley et al., 2021; Fan et al., 2020; Kreuter et al., 2020). The Delphi Group at CMU administers the U.S. COVID-19 Trends and Impact Survey (Delphi US CTIS) (Delphi Group, CMU, 2020), which collects approximately 50,000 responses per day from Facebook users in the United States. This survey includes self-reported information on symptoms, testing, vaccination, behaviors, and other COVID-19-related topics. Access to the data was obtained through an agreement with CMU, UMD, and Facebook, as described in the Ethical Declaration section. The full set of survey questions in the U.S. is available at (Delphi Group, Carnegie Mellon University, [n. d.]).
@@
Specifically, we computed daily aggregated information considering questions about age, education, source of information, and infection-induced fear—a summary of the selected questions and their corresponding response options is provided in Table 1. Importantly, these daily aggregates were calculated using the weights provided in the Delphi US CTIS microdata files. These weights were designed to address the differences between Facebook users and the overall US population, as well as to correct for biases associated with non-response (Salomon et al., 2021). For clarity and simplicity, throughout the remainder of the paper, we refer to responses to D2 as Age, D8 as Education, I5 as Source, and G1 as Fear. For Source, users could select multiple options, whereas for Age, Education, and Fear responses are mutually exclusive. Therefore, we accounted for all possible combinations of the nine response options in Source, totaling 256—also comprising the None of the above response, which is used in mutual exclusion with all the other possible answers.
In the remainder of the paper, we will focus on the relationships between Age, Education, and Source with respect to COVID-19–induced Fear, employing the notation presented in Table 2. Our analysis covers the period from May 2021 to June 2022 across all U.S. states, utilizing data from digital platforms such as Facebook, which combine online activity with self-reported information.
@@
(1) $\hat{\sigma}_{s}(v)=\frac{\sigma_{s}(v)}{\sum_{i\in\mathcal{S}}\sigma_{i}(v)},$
@@
(2) $\hat{\sigma}_{s^{\prime}}(v)=\sum_{\forall u\in\mathcal{S}:s^{\prime}\in u}\hat{\sigma}_{u}(v),$
(3) $\hat{\varphi}_{s}(v)=\sum_{f\in\mathcal{F}}w_{f}\cdot\hat{\varphi}_{s,\ f}(v),$
@@
$\sigma_{s}$ Avg. number of respondents selecting the combination $s\in\mathcal{S}$
$\sigma_{s^{\prime}}$ Avg. number of respondents selecting singleton $s^{\prime}\in\mathcal{S}$
$\varphi_{s,\ f}$ Avg. number of respondents who self-reported a fear level $f\in\mathcal{F}$, selecting the combination $s\in\mathcal{S}$
$\varphi_{s}$ Avg. fear score among respondents selecting the combination $s\in\mathcal{S}$
$\phi_{s^{\prime}}$ Estimate of avg. fear score for singletons $s^{\prime}\in\mathcal{S}$
@@
Table 2. Notation used in the paper.
@@
(4) $\min_{w_{s^{\prime}}|s}\sum_{t\in T}\left(\hat{\varphi}_{s}(v,\ t)-\sum_{s^{\prime}\in s}w_{s^{\prime}|s}\cdot\hat{\varphi}_{s^{\prime}}(v,\ t)\right)^{2},$
@@
(5) $\phi_{s^{\prime}}(v)=\sum_{s\in\mathcal{S}}w_{{s^{\prime}}|s}\cdot\hat{\varphi}_{s^{\prime}}(v)$
4. Statistical analysis of fear

### v2

@@ Abstract
The COVID-19 pandemic triggered not only a global health crisis but also an infodemic, an overload of information from diverse sources influencing public perception and emotional responses. In this context, fear emerged as a central emotional reaction, shaped by both media exposure and demographic factors. In this study, we analyzed the relationship between individuals’ self-reported levels of fear about COVID-19 and the information sources they rely on, across nine source categories, including medical experts, government institutions, media, and personal networks. In particular, we defined a score that ranks fear levels based on self-reported concerns about the pandemic, collected through the Delphi CTIS survey in the United States between May 2021 and June 2022. We found that both fear levels and information source usage closely follow COVID-19 infection trends, exhibit strong correlations within each group (fear levels across sources are strongly correlated, as are patterns of source usage), and vary significantly across demographic groups, particularly by age and education. Applying causal inference methods, we found that among age, education, and information source, the latter is the most influential factor affecting individuals’ fear levels. We further quantified the impact of different information sources on fear by estimating the average treatment effect, indicating how each source alters fear relative to a control. Furthermore, we demonstrated that information source preferences can reliably match the political orientation of U.S. states. These findings highlight the importance of information ecosystem dynamics in shaping emotional and behavioral responses during large-scale crises.
Introduction
@@
The COVID-19 pandemic represented a global health crisis of unprecedented scale and complexity. At the same time, it generated an infodemic—a rapid and overwhelming surge of information from a wide range of sources (Cinelli and others 2020). As individuals sought clarity and guidance, they turned to news media, social platforms, government announcements, and public health organizations. The credibility, consistency, and framing of these sources played a critical role in shaping public understanding of the threat. Additionally, they influenced emotional responses across populations (Jones et al. 2021; Lu and others 2021).
@@
In this context, fear emerged as one of the most prominent psychological responses to the pandemic (Coelho et al. 2020). This emotion was not driven solely by the objective health threat, but also by the volume, tone, and perceived trustworthiness of the information consumed. The impact of these information sources on fear varied across different demographics. Differences in age, education, gender, and socioeconomic status significantly shaped how individuals selected, interpreted, and trusted various sources (Nino et al. 2021). These disparities highlight the necessity of accounting for both the nature of the information disseminated and the characteristics of the recipients when evaluating the psychological impact of information exposure.
@@
Figure 1: Schematic representation of our study.
Contributions
@@
Our main contributions are as follows—summarized in Figure 1. First, using statistical tests and correlation analyses on U.S. data, we quantify the influence of different information sources on individuals’ COVID-19–related fear levels, and similarly assess differences in information source usage across the population. Second, we investigate how demographic factors—specifically age and education—not only affect emotional responses but also significantly shape information-seeking behavior and trust, revealing notable patterns across population groups. Third, through a causal inference model, we compare the relative impact of age, education, and information source on fear, finding that information source is the most influential driver. We then estimate the average treatment effect (ATE) to quantify how each source alters fear compared with a control source. Finally, building on this, we apply K-means, spectral, and hierarchical clustering to group U.S. states based on their residents’ reliance on different types of information. Remarkably, the resulting clusters exhibit a strong alignment with state-level political orientation—specifically, whether a state tends to lean Democratic or Republican. Collectively, our findings underscore the critical role of information sources in shaping population emotional responses to large-scale crises. They also highlight the dangers of misinformation—particularly when disseminated by institutional or authoritative entities—which can distort public perception, exacerbate fear, and undermine trust in democratic institutions.
@@
This article is organized as follows. The Related work section reviews prior research on fake news, misinformation, and COVID-19–related fear, highlighting key findings and existing gaps. The Data section describes our data sources and the construction of the indices used in the analysis. The Methods section presents the methodology, including correlation analysis, causal inference, and clustering. The Results section examines variation in fear levels across information sources and demographic groups, as well as variation in information source usage across demographics, and identifies patterns of information reliance across the U.S. Finally, the Conclusion section summarizes the main findings and discusses their broader implications.
Related work
A growing body of research has examined the spread of fake news and misinformation during the COVID-19 pandemic and their psychological consequences—particularly the prevalence and dynamics of fear. This fear has been fueled by false or low-credibility information presented as legitimate news which has also undermined trust in vaccines and preventive measures and contributed to a global infodemic (Caceres and others 2022).
Evidence shows that information sources strongly shape risk perceptions and emotional responses. In Portugal, scientific sources had greater effects on perceived personal and societal risk than general media, and trust in science heightened perceived vulnerability, while trust in social media reduced it (Entradas 2022). In the U.S., socioeconomic and political divisions shaped information-seeking behaviors: individuals with higher education and digital skills accessed a more diverse range of sources, whereas reliance on conservative media was associated with lower levels of factual knowledge (Reisdorf and others 2021). Additional analyses of social media behavior show that online discourse around COVID-19 was heavily politicized: an investigation of Facebook posts from users who engaged with anti-government and COVID-denialist narratives before later dying from the virus found that harmful themes were largely propagated within a right-wing political and media ecosystem, with platforms applying only soft moderation measures (Habib and Nithyanand 2023). This underscores both the politicization of COVID-19 information and the role of platform governance in shaping exposure to dangerous narratives.
False information spreads far more rapidly than accurate content on social platforms, largely propagated by superspreaders in ideologically aligned networks. While fact-checking demonstrates limited durability, psychological inoculation strategies employing gamified prebunking show enhanced efficacy in building cognitive resistance (Van Der Linden 2022). These dynamics amplify infectious disease transmission through behavioral mechanisms by promoting harmful practices and undermining prevention. Misinformation persists longer and spreads more readily than factual information, substantially increasing infection rates (Bernardin and Perez-Acle 2025). Additionally, the impact on vaccination intentions is especially pronounced; randomized controlled trials demonstrate that exposure to vaccine misinformation reduces the intent to vaccinate by approximately 6 percentage points relative to factual information, with greater effects for scientifically sounding misinformation and differential impacts across sociodemographic groups (Loomba and others 2021). Research also shows that misinformation about COVID-19 vaccines is driven by coordinated conspiracy-oriented communities that amplify anti-vaccine narratives and distort scientific evidence, contributing to polarized uptake patterns (Sharma et al. 2022). Automated approaches to misinformation monitoring also support these findings: neural stance-classification models using graph attention networks can reliably identify whether social media posts adopt or reject COVID-19 misinformation, revealing consistent patterns in the emotional, lexical, and thematic structure of misinformation uptake (Weinzierl et al. 2021). Meanwhile, even a single trusted correction can reduce belief in misinformation (Seo et al. 2022; Miyazaki and others 2023), emphasizing the importance of timely and credible debunking.
Social media commentary on COVID-19 was largely negative, with anger dominating globally and fear peaking early in the pandemic, later rising only during variant or mortality news. Emotional expression varied by region, with the U.S. showing more anger, African outlets more sadness, and Indian sources more fear. In particular, fear emerges as a central psychological response shaped more by subjective perceptions and social transmission than by epidemiological indicators. Social media analyses reveal that fear expression peaks during early outbreak phases and shows substantial gender differences, with women expressing fear more frequently than men in response to disease-related topics. The temporal dynamics of fear demonstrate marked fluctuations independent of actual disease severity, with daily shares of fear posts significantly elevated during peak COVID-19 periods compared to baseline 2019 levels (Chai and others 2022). Cross-national longitudinal research finds that roughly 23% of fear variance is explained by micro-level psychological factors—worries about resource shortages, perceived disease vulnerability, germ aversion, and infections in close social circles—while macro-level indicators like case counts, mortality, and restrictions contribute minimally. Individual traits such as age, attachment anxiety, and prior health concerns further shape fear responses, revealing heterogeneous susceptibility and a disconnection between subjective threat perception and objective epidemiological data (Eder and others 2021). Agent-based models with associative learning reveal that falling fear can trigger new disease waves by reducing protective behavior, which strongly shapes epidemic trajectories. Media campaigns promoting protective actions have minimal impact, suggesting interventions must target the psychological drivers of fear and compliance, not just factual disease information (Retzlaff and others 2022).
Moreover, early population-level fear modestly predicted later vaccination willingness but was insufficient alone to explain vaccine hesitancy, which involves multiple psychological and risk-perception factors. Demographically, females, healthcare workers, lower-educated, and married individuals showed higher fear, highlighting vulnerable groups needing targeted interventions that address fear alongside vaccine-related beliefs (Oliveira and others 2023; Mertens et al. 2022; Doshi and others 2021). Finally, causal inference reveals bidirectional relationships between media exposure, interpersonal communication, and fear of COVID-19. Both media frequency and perceived danger in conversations predict fear, while elevated fear motivates increased media and communication seeking. Counterarguing proves most effective at reducing fear and encourages sustained use, with individuals cycling from avoidance to counterarguing to reappraisal as the pandemic progresses. These adaptive emotion-regulation patterns suggest that targeted interventions addressing maladaptive regulatory pathways prove more effective than uniform communication approaches (Li 2022).
Overall, misinformation and fear jointly influence pandemic behavior through interconnected mechanisms. Misinformation spreads rapidly, reduces vaccination intentions, and undermines protective behaviors, while fear—driven largely by subjective perceptions, information exposure, and social context—varies across demographic groups and feeds back into information-seeking patterns. Declining fear can weaken compliance and contribute to renewed outbreaks.
@@
(I5) In the past 7 days, from which of the following sources have you received news and information about COVID-19? (1) Doctors, (2) Scientists, (3) CDC, (4) Government, (5) Politicians, (6) Journalists, (7) Family and friends, (8) Religious leaders, (9) None of the above (other sources not listed here).
@@
Table 1: Selected questions and responses from the Delphi US CTIS survey, with labels as in the original survey.
@@
Despite extensive research on misinformation and fear during the COVID-19 pandemic, key gaps remain. The influence of different information sources on fear across demographic groups remains poorly characterized. Additionally, geographic and political influences remain underexplored. Addressing these gaps is critical for designing interventions that effectively mitigate misinformation, target heterogeneous fear responses, and support vulnerable populations. Our study addresses these gaps by employing advanced statistical and causal inference methods to evaluate how demographics and sources of information jointly shape fear responses. Moreover, by clustering U.S. states according to patterns of information source usage and linking these clusters to political orientation, we offer a comprehensive, integrated perspective on the social and contextual determinants of COVID-19 fear.
Data
@@
Throughout this work, we relied on two sources of COVID-19 data: official surveillance data from the COVID-19 Data Hub (Guidotti and Ardia 2020; Guidotti 2022) and survey-based data from the Delphi US CTIS (Delphi Group, CMU 2020), which contains information on demographics, the types of information sources used by individuals, and self-reported levels of fear. All datasets used in this study are open-source and publicly available. Aggregated data from the Delphi US CTIS survey can be accessed as described in the Availability of materials and data section.
@@
We primarily used the COVID-19 Data Hub (Guidotti and Ardia 2020; Guidotti 2022) as our source of surveillance data. Specifically, we used daily new fatalities and confirmed cases—computed from cumulative counts—as well as the total U.S. population. These indicators were used to estimate the daily number of official infections via Sybil (Baccega et al. 2024). Specifically, this data allows us to reconstruct the dynamics of a Susceptible-Infected-Recovered-Deceased-Susceptible (SIRDS) compartmental model with reinfections over the study period. In this model, the infected compartment represents the estimate of the official number of infections (Álvarez and others 2021; Astley and others 2021)—we assumed an average recovery time of 14 days to compensate for the lack of recovery data, which are largely unavailable. While the Data Hub provides additional information, such as vaccination rates and hospital statistics, these were not relevant to our study.
@@
Since the spring of 2020, Facebook has collaborated with both Carnegie Mellon University (CMU) and the University of Maryland (UMD) through its Data for Good program (Facebook 2020) to promote large-scale, daily surveys developed by these institutions (Astley and others 2021; Fan and others 2020). The Delphi Group at CMU administers the U.S. COVID-19 Trends and Impact Survey (Delphi Group, CMU 2020)—Delphi US CTIS—, which collects approximately 50,000 responses per day from Facebook users in the United States. This survey includes self-reported information on symptoms, testing, vaccination, behaviors, and other COVID-19-related topics, and was used to estimate prevalence trends across U.S. states, accounting for direct and indirect reporting (Srivastava et al. 2024). Access to the data was obtained through an agreement with CMU, UMD, and Facebook, as described in the Ethical Declaration section. The full set of survey questions in the U.S. is available at (Delphi Group, Carnegie Mellon University 2020).
@@
Specifically, we computed daily aggregated information considering questions about age, education, source of information, and infection-induced fear—a summary of the selected questions and their corresponding response options is provided in Table 1. Importantly, these daily aggregates were calculated using the weights provided in the Delphi US CTIS microdata files. These weights were designed to address the differences between Facebook users and the overall US population, as well as to correct for biases associated with non-response (Salomon and others 2021). For clarity and simplicity, throughout the remainder of the paper, we refer to responses to D2 as Age, D8 as Education, I5 as Source, and G1 as Fear. For Source, users could select multiple options, whereas for Age, Education, and Fear responses are mutually exclusive. Therefore, we accounted for all possible combinations of the nine response options in Source, totaling 256—also comprising the None of the above response, which is used in mutual exclusion with all the other possible answers. We will focus on the relationships between Age, Education, and Source with respect to COVID-19–induced Fear, employing the notation presented in Table 2.
Our analysis covers the period from May 2021 to June 2022 across all U.S. states, which combines online activity with self-reported information. While these data sources are inherently biased toward individuals with consistent internet access—specifically on active Facebook users—and may underrepresent certain socio-demographic groups or regions with limited connectivity, these biases are less pronounced in U.S. states where internet access is widespread—hence, offering valuable insights into broader behavioral trends.
@@
$\hat{\sigma}_{s}(v)=\frac{\sigma_{s}(v)}{\sum_{i\in\mathcal{S}}\sigma_{i}(v)},$ (1)
@@
$\sigma_{s}$ Average number of respondents selecting the combination $s\in\mathcal{S}$
$\sigma_{s^{\prime}}$ Average number of respondents selecting singleton $s^{\prime}\in\mathcal{S}$
$\varphi_{s,\ f}$ Average number of respondents who self-reported a fear level $f\in\mathcal{F}$, selecting the combination $s\in\mathcal{S}$
$\varphi_{s}$ Average fear score among respondents selecting the combination $s\in\mathcal{S}$
$\phi_{s^{\prime}}$ Estimate of average fear score for singletons $s^{\prime}\in\mathcal{S}$
@@
Table 2: Notation used in the paper.
$\hat{\sigma}_{s^{\prime}}(v)=\sum_{\forall u\in\mathcal{S}:s^{\prime}\in u}\hat{\sigma}_{u}(v),$ (2)
$\hat{\varphi}_{s}(v)=\sum_{f\in\mathcal{F}}w_{f}\cdot\hat{\varphi}_{s,\ f}(v),$ (3)
@@
$\min_{w_{s^{\prime}}|s}\sum_{t\in T}\left(\hat{\varphi}_{s}(v,\ t)-\sum_{s^{\prime}\in s}w_{s^{\prime}|s}\cdot\hat{\varphi}_{s^{\prime}}(v,\ t)\right)^{2},$ (4)
@@
$\phi_{s^{\prime}}(v)=\sum_{s\in\mathcal{S}}w_{{s^{\prime}}|s}\cdot\hat{\varphi}_{s^{\prime}}(v)$ (5)
Only weights from models with statistically significant coefficients—determined via F-tests—are retained to ensure robust attribution.
Methods
This section describes the primary methodology used in our study, as illustrated in Figure 1.
Statistical analysis on fear. Using U.S. data, we assess how different information sources affect COVID-19–related fear through statistical tests—Shapiro–Wilk (Shapiro and Wilk 1965), Kolmogorov–Smirnov (Massey Jr 1951) and Mann–Whitney U (Nachar and others 2008) tests—and Spearman correlation analyses (Spearman 1961). We further examine how demographic factors—specifically age and education—shape emotional responses, thereby revealing distinct patterns across population groups. Specifically, the Shapiro–Wilk test was used to assess whether a distribution follows normality, whereas the Kolmogorov–Smirnov test was applied to compare two distributions, evaluating differences across their full range, including shape, dispersion, and central tendency. In contrast, the Mann–Whitney U test was employed to detect differences in central tendency between two distributions. The Spearman correlation coefficient was used to quantify the strength and direction of monotonic relationships, allowing us to examine how fear levels co-vary with different information sources.
Causal inference. Using causal inference, we evaluate the relative contributions of age, education, and information source to COVID-19–related fear, finding that information source is the most influential factor. In particular, causal inference aims to uncover cause-and-effect relationships rather than mere statistical associations. It employs a range of statistical methods and study designs—such as randomized controlled trials and quasi-experimental approaches—to estimate the effect of an exposure or intervention on an outcome while accounting for potential confounders. Randomized controlled trials remain the gold standard for establishing causality, as randomization helps balance confounding factors across groups. However, these trials are often infeasible or ethically problematic. In such cases, causal inference methods can be applied to observational data, provided key assumptions hold—consistency, positivity, and exchangeability. Moreover, we estimate the average treatment effect (ATE) of different information sources, where the ATE—one of the primary metrics in causal inference—quantifies the expected difference in outcomes between a control treatment (in our case, None of the above) and each alternative information source, thereby reflecting the causal impact of the intervention at the population level.
Figure 2: Causal inference model used.
To define the causal inference model—shown in Figure 2—, we use the DoWhy package (Sharma and Kiciman 2020; Blöbaum and others 2024), a Python library that integrates graphical causal models with statistical methods to specify assumptions, identify estimands, estimate causal effects, and perform robustness checks, including refutation tests. DoWhy was mainly used to estimate the strength of causal relationships (arrow strengths) by fitting a Probabilistic Causal Model (PCM) to the data. For effect estimation, however, we used Ordinary Least Squares (OLS) regression models, adjusting for confounders (Age and Education) identified in the causal graph, to obtain ATE estimates. DoWhy’s refutation API was applied to validate model robustness through tests like placebo treatments and subset analyses, ensuring the reliability and interpretability of our OLS-based causal estimates from observational data.
In detail, the positivity assumption was verified by confirming that all information sources appear across all Age and Education strata, ensuring each source has a positive probability of selection. Each source was treated as a binary treatment (selected or not on a given date), assuming consistent definitions, no meaningful heterogeneity in usage, and additive effects. While the intensity and context of source use may vary (e.g., passive exposure vs. active seeking), such distinctions are not captured in the survey. The analysis further assumes that age and education suffice to control for confounding, though unmeasured factors—such as political ideology, institutional trust, socioeconomic status, and personal traits—may influence results. Consequently, estimated causal effects reflect associations conditional on measured confounders, and causal interpretation relies on the unverifiable assumption of no relevant unmeasured confounding. Sensitivity analyses for unmeasured confounding are beyond the scope of this study.
Geographical distribution of sources. Finally, we statistically assess how people use different information sources using the previously described tests and Spearman correlation analyses, while also examining how demographic factors influence information source usage, thereby revealing distinct patterns across population groups. Next, we apply K-means (Hartigan et al. 1979), spectral (Ng et al. 2001), and hierarchical clustering (Johnson 1967) to group U.S. states based on their residents’ reliance on different types of information. Remarkably, the resulting clusters exhibit a strong alignment with state-level political orientation—specifically, whether a state tends to lean Democratic or Republican. In detail, K-means partitions data into $k$ clusters by iteratively assigning points to the nearest centroid, while spectral clustering uses graph Laplacian eigenvectors to capture non-convex structures before applying a partitioning method such as K-means. Hierarchical clustering builds a dendrogram through agglomerative or divisive steps, allowing flexible cluster counts but at a higher computational cost. Cluster quality is commonly assessed using the Silhouette score, which quantifies how well points fit within their assigned clusters.
Results